# Using Fiber Optic Bundles to Miniaturize Vision-Based Tactile Sensors

### v2

@@ Appendix C Silicone Preparation for Hardness Experiments
Roberto Calandra is a Full (W3) Professor at the Technische Universität Dresden, and at the Centre for Tactile Internet with Human-in-the-Loop (CeTI). Previously, he founded at Meta AI (formerly Facebook AI Research) the Robotic Lab in Menlo Park. Prior to that, he was a Postdoctoral Scholar at the University of California, Berkeley (US) in the Berkeley Artificial Intelligence Research (BAIR) Lab. His education includes a Ph.D. from TU Darmstadt (Germany), a M.Sc. in Machine Learning and Data Mining from the Aalto university (Finland), and a B.Sc. in Computer Science from the Università degli studi di Palermo (Italy). His scientific interests are broadly at the conjunction of Robotics and Machine Learning, with the goal of making robots more intelligent and useful in the real world. Among his contributions is the design and commercialization of DIGIT – the first commercially available high-resolution compact tactile sensor. Roberto served as Program Chair for AISTATS 2020, as Guest Editor for the JMLR Special Issue on Bayesian Optimization, and has previously co-organized over 16 international workshops (including at NeurIPS, ICML, ICLR, ICRA, IROS, RSS). |
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_di-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_dugonjic-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_fu-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_wu-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_mercado-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_sawyer-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_most-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_kammerer-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_speidel-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_fan-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_sonn-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_cutkosky.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/photo_lambeta-min.jpg)
![[Uncaptioned image]](extracted/5464226/fig/biography/roberto_calandra.jpg)

### v3

@@ Appendix I Additional Results From Studying Image Sequence Length
Roberto Calandra is a Full (W3) Professor at the Technische Universität Dresden. Previously, he founded at Meta AI (formerly Facebook AI Research) the Robotic Lab in Menlo Park. Prior to that, he was a Postdoctoral Scholar at the University of California, Berkeley (US). His education includes a Ph.D. from TU Darmstadt (Germany), a M.Sc. in Machine Learning and Data Mining from the Aalto university (Finland), and a B.Sc. in Computer Science from the Università degli studi di Palermo (Italy). His scientific interests are broadly at the conjunction of Robotics, Touch Sensing, and Machine Learning. In 2024, he received the IEEE Early Academic Career Award in Robotics and Automation. |
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_di-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_dugonjic-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_fu-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_wu-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_mercado-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_sawyer-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_most-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_kammerer-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_speidel-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_fan-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_sonn-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_cutkosky.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/photo_lambeta-min.jpg)
![[Uncaptioned image]](extracted/5943887/fig/biography/roberto_calandra.jpg)